A insulator sphere had a charge volume density 2.4 µC/m3 and it’s radius is
9.6 cm.What is the electric field at 5cm from it’s center ?
1
Expert's answer
2020-10-30T13:13:07-0400

The electric field inside a charged sphere is given by:


"E = \dfrac{4}{3}\rho kr"

where "\rho = 2.4\times 10^{-6}C/m^3" is the charge volume density, "k = 9\times 10^9N\cdot m^2/C^2" is the constant. Thus, if "r = 5cm = 0.05m", obtain:

"E = \dfrac{4}{3}\rho kr = \dfrac{4}{3} 2.4\times 10^{-6}\cdot 9\times 10^9\cdot 0.05 =1.44\space V/m"

Answer. 1.44 V/m.